If 3.0x105 J of work was done moving a box 10.0 m, calculate the force applied to the box over that distance.
1
Expert's answer
2020-05-19T10:45:26-0400

The work needed to move an object by the constant force "F" by displacement "s" is "A = F \cdot s", from where "F = \frac{A}{s} = \frac{3.0 \cdot 10^5 J}{10 m} = 3 \cdot 10^4 N = 30kN".